OpenStudy (anonymous):

use the distributive property to simplify the expression 2x-3[5x-2(6-x)]

OpenStudy (espex):

Distribute your -2, then your -3, then combine like terms and simplify.

OpenStudy (anonymous):

i think it's 36-19x

OpenStudy (comm.dan):

After you are done getting your answer, tell us what you get. @melanie083194

OpenStudy (anonymous):

First 'Distribute' the -2 to each term in (6 - x). After you've done that and simplified, distribute 2x to every term in the big bracket and -3 to everything in the big bracket. \[\bf (a+b)(c+d)=a(c+d)+b(c+d)=ac+ad+bc+bd\]

OpenStudy (anonymous):

@Comm.Dan ok i took the negative 3 and multiplied it with the 5x and got -15x and then the -3 to the -2 and got +6 then the -2 to the 6 in the () and got -12 and then negative to the -x and got -2x

OpenStudy (comm.dan):

Could you please do it in the equation tab? @melanie083194

OpenStudy (espex):

@melanie083194 you need to distribute the -2 first, THEN you can distribute the -3 in.

OpenStudy (anonymous):

What is the equation of the line passing through the points (–2, 3) and (1, 4)?